# Uni-3DAR: Unified 3D Generation and Understanding via Autoregression on Compressed Spatial Tokens

### v2

@@ 3.2 Crystal Generation
We benchmark Uni-3DAR against established methods, including FTCP [43], G-SchNet [44], P-G-SchNet [44], CDVAE [37], DiffCSP [38], and FlowMM [39]. Additionally, we evaluate Uni-3DAR against the recent UniGenX [45] for the CSP task. For PXRD-guided CSP, we compare Uni-3DAR with PXRDGEN [46], a model tailored for this task.
In Uni-3DAR, we use a 12-layer model with a 768-dimensional embedding for de novo crystal generation, while a larger 24-layer model with a 1024-dimensional embedding is employed for CSP and PXRD-guided CSP tasks. All models are trained for 400k steps with a batch size of 64 and a peak learning rate of 3e-4. For chemical composition conditioning, we prepend a token derived from a multi-hot atom-type vector. PXRD data, spanning angles from 0° to 120°, is converted into a 1200-dimensional vector with a 0.1° resolution, evenly divided into four segments, each represented by a conditional token. As a result, PXRD-guided CSP utilizes a total of five conditional tokens (one for composition and four for PXRD signals). The autoregressive nature of Uni-3DAR enables seamless integration of these conditional tokens, eliminating the need for additional encoders required by previous methods [46, 47].
@@
Table 4 demonstrates Uni-3DAR’s performance in PXRD-guided CSP on the MP-20 dataset, benchmarked against PXRDGEN [46]. Uni-3DAR substantially outperforms PXRDGEN, elevating the match rate from 68.68% to 75.08% while drastically reducing the RMSE from 0.0707 to 0.0276—a 256% relative improvement. This significant RMSE reduction underscores Uni-3DAR’s exceptional ability to generate crystal structures that precisely correspond to experimental PXRD patterns. Collectively, these results underscore the superior capability of Uni-3DAR in harnessing diffraction constraints to reliably predict crystal structures.
3.3 Protein Pocket Prediction
@@
We follow previous studies [48] and employ a binding site dataset constructed from the CASF-2016 core set [49], PDBBind v2020 refined set [50], and MOAD [51]. The dataset consists of 23k training samples, 5k validation samples, and five test sets of roughly 1k samples each. Model performance is assessed using the Intersection-over-Union (IoU) metric, consistent with previous evaluations [48].
@@
We benchmark Uni-3DAR against established methods. Our comparisons include non-pretrained approaches (e.g., FPocket [52], SiteHound [53], etc.) and pretrained models (e.g., ESM2_150M [54], GearNet [55], Siamdiff [56], and Vabs-Net [48]). In line with prior works [48], we pretrain Uni-3DAR on approximately 1.3 million protein structures before fine-tuning it on the binding site dataset. Unlike Vabs-Net, which employs full-atom representations, our experiments are restricted to $\alpha$-carbon atoms to facilitate direct comparisons.
@@
Results
As summarized in Table 5, Uni-3DAR achieves state-of-the-art performance compared to other baselines. Importantly, Uni-3DAR reaches comparable or superior results despite using only 3D protein structures, whereas methods like Vabs-Net additionally leverage ESM embeddings and Solvent Accessible Surface Area. These findings underscore Uni-3DAR’s strong capacity for accurately understanding protein 3D structures and effectively performing atom-level tasks.
@@ Results
Method pretrained B277$\uparrow$ DT198$\uparrow$ ASTEX85$\uparrow$ CHEN251$\uparrow$ COACH420$\uparrow$
FPocket [52] $\times$ 31.5 23.2 34.1 25.4 30.0
SiteHound [53] $\times$ 36.4 23.1 38.9 29.4 34.9
MetaPocket2 [57] $\times$ 37.3 25.8 37.5 32.8 37.7
DeepSite [58] $\times$ 34.0 29.1 37.4 27.4 33.9
P2Rank [59] $\times$ 49.8 38.6 47.4 56.5 45.3
ESM2_150M [54] $\surd$ 19.6 16.6 20.5 18.9 22.0
GearNet [55] $\surd$ 39.9 35.8 41.0 36.4 41.3
Siamdiff [56] $\surd$ 37.7 31.0 40.7 35.3 40.3
Vabs-Net [48] $\surd$ - - - - 56.3
@@
Uni-Dock [61] 32.51$\pm$0.39 50.69$\pm$0.59 1.89$\pm$0.04 47.11$\pm$0.22 67.03$\pm$0.94 1.10$\pm$0.02
Glide SP [62] 17.36$\pm$0.00 44.63$\pm$0.00 2.27$\pm$0.00 31.13$\pm$0.00 60.06$\pm$0.00 1.54$\pm$0.00
GNINA [63] 21.12$\pm$0.26 43.62$\pm$1.06 2.45$\pm$0.07 28.47$\pm$0.57 58.13$\pm$0.81 1.65$\pm$0.02
SMINA [64] 18.73$\pm$0.00 31.68$\pm$0.00 3.99$\pm$0.00 28.47$\pm$0.56 48.48$\pm$0.00 2.07$\pm$0.00
Vina [65] 18.32$\pm$0.02 36.64$\pm$0.05 3.42$\pm$0.01 24.79$\pm$0.00 50.96$\pm$0.00 1.87$\pm$0.01
EquiBind [66] / 5.5$\pm$1.2 6.2$\pm$0.3 / / /
TANKBind [67] 2.66$\pm$0.26 18.18$\pm$0.60 4.2$\pm$0.05 4.13$\pm$0.0 20.39$\pm$0.45 3.5$\pm$0.04
E3Bind [68] / 25.6 7.2 / / /
KarmaDock [69] / 56.2 / / / /
DiffDock(Pocket) [70] / 51.8 2.0 / 60.7 1.9
DiffDock [70] 15.15 36.09 3.35 21.76 43.52 2.46
DiffDock-L [71] 19.07$\pm$0.57 40.74$\pm$1.25 2.88$\pm$0.18 21.95$\pm$0.39 48.15$\pm$0.91 2.05$\pm$0.04
SurfDock [60] 40.96$\pm$0.34 68.41$\pm$0.26 1.18$\pm$0.00 54.18$\pm$0.13 75.11$\pm$0.13 0.94$\pm$0.00
Uni-3DAR 44.75$\pm$2.63 69.06$\pm$0.75 1.08$\pm$0.04 56.35$\pm$1.99 72.38$\pm$0.73 0.76$\pm$0.02
3.4 Molecular Docking
@@
We evaluate Uni-3DAR against 13 baselines, including five classical docking software tools and eight deep learning–based methods. Most existing deep learning approaches rely on complex featurizations, such as using protein language model embeddings (e.g., from ESM2 [72]). To simplify and unify the molecular tasks, we omit these complicated features in Uni-3DAR and instead use only atom types and coordinates. We also adopt a full-atom representation of the protein pocket to enhance expressive power. We frame docking as an autoregressive generation task by embedding both the pocket and the RDKit conformation as two frames, concatenating them into a single input sequence, and training the model to generate the docked molecule conformation as a new frame sequence. For further simplicity, we do not impose constraints such as matching the number and types of atoms in the output frame to those of the input molecule. Also, we do not train a separate scoring model for pose ranking. Instead, we use the cumulative probability derived from autoregressive generation to score each generated pose. We train Uni-3DAR for 300k steps (approximately 300 epochs) with a batch size of 16. The learning rate schedule follows the same configuration as the experiments detailed in section 3.1. The training is completed in approximately one day on 4 NVIDIA A100 GPUs.
@@ Results
Experimental results are summarized in Table 6. Uni-3DAR outperforms to the state-of-the-art method, SurfDock, demonstrating similar percentages of poses with RMSD below 1Å and 2Å. Notably, Uni-3DAR excels in generating higher-quality poses, reflected by its lower median RMSD values. However, Uni-3DAR exhibits slightly inferior performance in selecting Top-5 poses for challenging cases, as evidenced by a lower percentage of poses with RMSD below 2Å in the Top5-RMSD evaluation (72.38% vs. 75.11% for SurfDock). This gap may arise because the scoring module in Uni-3DAR has not been explicitly trained and is only exposed to ground-truth conformations during the training phase. Addressing this limitation by training a dedicated scoring module could potentially enhance its selection performance. Moreover, since Uni-3DAR avoids complex feature engineering, its docking accuracy might further benefit from multitask learning strategies, emphasizing the promise of a unified foundational model for molecular applications.
3.5 Molecular Property Prediction via Pretraining
@@
We utilize the same pretraining dataset as employed by Uni-Mol [8] and SpaceFormer [73], comprising approximately 19 million molecules. For downstream evaluations, we follow the datasets and evaluation settings used by the state-of-the-art SpaceFormer [73]. These include a 20K dataset predicting electronic properties (HOMO, LUMO, GAP), a 21K dataset targeting energy properties (E1-CC2, E2-CC2, f1-CC2, f2-CC2), and an 8K dataset predicting mechanical and electronic properties (Dipmom, aIP, and D3 Dispersion Corrections). Data splits align exactly with SpaceFormer’s methodology [73]. Performance across all tasks is measured using the Mean Absolute Error (MAE) metric.
@@
Our baselines encompass several prominent models, including Uni-Mol [8], Mol-AE [74], 3D Infomax [75], GROVER [76], GEM [77], and the most recent state-of-the-art method, SpaceFormer [73]. For pretraining, we use the proposed masked next-token prediction as pretraining task, training the model for 500k steps with a batch size of 128. The peak learning rate is set to 3e-4, incorporating a 10% linear warmup followed by cosine decay, requiring approximately 11.5 hours on 8 NVIDIA 4090 GPUs.
@@
Results
The experimental results summarized in Table 7 clearly illustrate the strong performance of Uni-3DAR. Uni-3DAR achieves top performance, ranking first in 4 out of the 10 tasks evaluated and placing within the top two positions in 8 out of 10 tasks. Notably, Uni-3DAR demonstrates comparable performance to the specialized SpaceFormer model, indicating that Uni-3DAR possesses robust representational capabilities competitive with state-of-the-art specialized representation models.
@@ Results
GROVER 0.0075 $\pm$ 2.0e-4 0.0086 $\pm$ 8.0e-4 0.0109 $\pm$ 1.4e-3 0.0101 $\pm$ 9.7e-4 0.0129 $\pm$ 4.6e-4 0.0219 $\pm$ 3.5e-4 0.0401 $\pm$ 1.2e-3 0.0752 $\pm$ 1.1e-3 0.1467 $\pm$ 1.5e-2 0.2516 $\pm$ 5.3e-2
GEM 0.0068 $\pm$ 7.0e-5 0.0080 $\pm$ 2.0e-5 0.0107 $\pm$ 1.9e-4 0.0090 $\pm$ 1.3e-4 0.0102 $\pm$ 2.3e-4 0.0170 $\pm$ 4.3e-4 0.0352 $\pm$ 5.4e-4 0.0289 $\pm$ 1.2e-3 0.0207 $\pm$ 2.6e-4 0.0077 $\pm$ 6.6e-4
3D Infomax 0.0065 $\pm$ 1.0e-5 0.0070 $\pm$ 1.0e-4 0.0095 $\pm$ 1.0e-4 0.0089 $\pm$ 2.0e-4 0.0091 $\pm$ 3.0e-4 0.0172 $\pm$ 4.0e-4 0.0364 $\pm$ 9.0e-4 0.0291 $\pm$ 1.7e-3 0.0526 $\pm$ 1.4e-4 0.2285 $\pm$ 7.5e-3
Uni-Mol 0.0052 $\pm$ 2.0e-5 0.0060 $\pm$ 6.0e-5 0.0081 $\pm$ 4.0e-5 0.0067 $\pm$ 4.0e-5 0.0080 $\pm$ 4.0e-5 0.0143 $\pm$ 2.0e-4 0.0309 $\pm$ 9.4e-4 0.0106 $\pm$ 3.1e-4 0.0095 $\pm$ 6.4e-4 0.0047 $\pm$ 5.6e-4
Mol-AE 0.0050 $\pm$ 8.0e-5 0.0057 $\pm$ 4.7e-4 0.0080 $\pm$ 8.0e-5 0.0070 $\pm$ 6.0e-5 0.0080 $\pm$ 4.0e-5 0.0140 $\pm$ 4.0e-5 0.0307 $\pm$ 1.3e-3 0.0113 $\pm$ 4.7e-4 0.0103 $\pm$ 1.3e-4 0.0077 $\pm$ 1.3e-3
SpaceFormer 0.0042 $\pm$ 1.0e-5 0.0040 $\pm$ 2.0e-5 0.0064 $\pm$ 1.2e-4 0.0058 $\pm$ 8.0e-5 0.0074 $\pm$ 8.4e-5 0.0142 $\pm$ 3.7e-4 0.0294 $\pm$ 7.1e-4 0.0083 $\pm$ 5.0e-4 0.0090 $\pm$ 5.9e-4 0.0053 $\pm$ 1.2e-3
Uni-3DAR 0.0048 $\pm$ 2.1e-5 0.0044 $\pm$ 3.2e-5 0.0065 $\pm$ 8.8e-5 0.0056 $\pm$ 2.2e-5 0.0067 $\pm$ 2.0e-5 0.0134 $\pm$ 7.0e-5 0.0286 $\pm$ 1.6e-4 0.0114 $\pm$ 6.9e-4 0.0127 $\pm$ 1.1e-4 0.0052 $\pm$ 3.2e-4
3.6 Polymer Property Prediction via Pretraining

### v3

@@ C.2 Crystal Generation
We benchmark Uni-3DAR against established methods, including FTCP [Ren et al., 2021], G-SchNet [Gebauer et al., 2019], P-G-SchNet [Gebauer et al., 2019], CDVAE [Xie et al., 2021], DiffCSP [Jiao et al., 2023], and FlowMM [Miller et al., 2024]. Additionally, we evaluate Uni-3DAR against the recent UniGenX [Zhang et al., 2025] for the CSP task. For PXRD-guided CSP, we compare Uni-3DAR with PXRDGEN [Li et al., 2024a], a model tailored for this task.
In Uni-3DAR, we use a 12-layer model with a 768-dimensional embedding for de novo crystal generation, while a larger 24-layer model with a 1024-dimensional embedding is employed for CSP and PXRD-guided CSP tasks. All models are trained for 400k steps with a batch size of 64 and a peak learning rate of 3e-4. For chemical composition conditioning, we prepend a token derived from a multi-hot atom-type vector. PXRD data, spanning angles from 0° to 120°, is converted into a 1200-dimensional vector with a 0.1° resolution, evenly divided into four segments, each represented by a conditional token. As a result, PXRD-guided CSP utilizes a total of five conditional tokens (one for composition and four for PXRD signals). The autoregressive nature of Uni-3DAR enables seamless integration of these conditional tokens, eliminating the need for additional encoders required by previous methods [Li et al., 2024a, Lai et al., 2025].
@@
Table 3 demonstrates Uni-3DAR’s performance in PXRD-guided CSP on the MP-20 dataset, benchmarked against PXRDGEN [Li et al., 2024a]. Uni-3DAR substantially outperforms PXRDGEN, elevating the match rate from 68.68% to 75.08% while drastically reducing the RMSE from 0.0707 to 0.0276—a 256% relative improvement. This significant RMSE reduction underscores Uni-3DAR’s exceptional ability to generate crystal structures that precisely correspond to experimental PXRD patterns. Collectively, these results underscore the superior capability of Uni-3DAR in harnessing diffraction constraints to reliably predict crystal structures.
C.3 Macroscopic 3D Object Generation
To demonstrate versatility beyond the microscopic realm of molecules and crystals, we further evaluate Uni-3DAR on unconditional macroscopic 3D object generation, a core task in 3D computer vision (3DCV). The goal is to synthesize realistic and diverse 3D shapes of everyday objects directly from the learned data distribution.
Dataset and Evaluation Protocol
Following common practice, we adopt three categories from ShapeNet [Chang et al., 2015]airplane, chair, and car. Each object is represented as a point cloud with 2,048 points uniformly sampled from the surface. In line with recent recommendations, we evaluate using 1-nearest-neighbor accuracy (1-NNA; lower is better) computed with both Chamfer Distance (CD) and Earth Mover’s Distance (EMD) [Yang et al., 2019, Vahdat et al., 2022]. Concretely, given a generated set $S_{g}$ and a reference set $S_{r}$, 1-NNA is the leave-one-out accuracy of a 1-NN classifier on $S_{g}\cup S_{r}$; if $S_{g}$ matches $S_{r}$ well, the classification accuracy approaches 50%. Compared with legacy metrics such as coverage (COV) and minimum matching distance (MMD), 1-NNA more directly captures distributional similarity while jointly reflecting both quality and diversity, and avoids several known failure modes of COV/MMD. We therefore report 1-NNA (with CD/EMD) as our primary metric throughout this section and in the main paper.
Baselines and Implementation Details
We benchmark Uni-3DAR against established point-cloud generative models, including r-GAN and l-GAN [Achlioptas et al., 2018], PointFlow [Yang et al., 2019], SoftFlow [Kim et al., 2020], SetVAE [Kim et al., 2021], DPF-Net [Klokov et al., 2020], diffusion-based methods DPM [Luo and Hu, 2021] and PVD [Zhou et al., 2021], and the recent LION [Vahdat et al., 2022]. To ensure clear and reproducible comparisons, we follow the PointFlow data protocol and training/test splits for the three categories.
For Uni-3DAR, the input 3D object is voxelized at $512\times 512\times 512$ resolution. We define fine-grained structural tokens as non-overlapping $16\times 16\times 16$ voxel patches and quantize each patch with a VQVAE codebook. (In our main text we summarize this as “each patch is quantized using VQVAE”; here we provide the fuller setup for completeness.) Unless otherwise specified, generation uses our single-frame sampling procedure, analogous to the molecular setting. For each ShapeNet category, the VQVAE is trained for 200 epochs; Uni-3DAR is then trained for 10,000 steps with batch size 64. On a single NVIDIA RTX 4090, training per category requires approximately 10 hours for the VQVAE and 2 hours for Uni-3DAR.
Results
Table 4 (main paper) summarizes unconditional generation under the 1-NNA protocol. Uni-3DAR achieves the lowest (best) 1-NNA in all six category–metric pairs (Airplane/Chair/Car $\times$ CD/EMD), outperforming strong diffusion and flow-based baselines. In particular, Uni-3DAR consistently improves over LION—the strongest baseline in our comparison—by small but systematic margins: Airplane (CD: 67.35 vs. 67.41; EMD: 61.09 vs. 61.23), Chair (CD: 53.11 vs. 53.70; EMD: 50.98 vs. 52.34), and Car (CD: 53.35 vs. 53.41; EMD: 50.89 vs. 51.14). Taken together, these results indicate that Uni-3DAR produces point-cloud distributions that are both high-quality and diverse, closely matching the real data according to a metric expressly designed to assess distributional similarity.
C.4 Protein Pocket Prediction
@@
We follow previous studies [Zhao et al., 2024] and employ a binding site dataset constructed from the CASF-2016 core set [Su et al., 2018], PDBBind v2020 refined set [pdb, 2025], and MOAD [Hu et al., 2005]. The dataset consists of 23k training samples, 5k validation samples, and five test sets of roughly 1k samples each. Model performance is assessed using the Intersection-over-Union (IoU) metric, consistent with previous evaluations [Zhao et al., 2024].
@@
We benchmark Uni-3DAR against established methods. Our comparisons include non-pretrained approaches (e.g., FPocket [Le Guilloux et al., 2009], SiteHound [Hernandez et al., 2009], etc.) and pretrained models (e.g., ESM2_150M [Lin et al., 2023b], GearNet [Zhang et al., 2022b], Siamdiff [Zhang et al., 2023d], and Vabs-Net [Zhao et al., 2024]). In line with prior works [Zhao et al., 2024], we pretrain Uni-3DAR on approximately 1.3 million protein structures before fine-tuning it on the binding site dataset. Unlike Vabs-Net, which employs full-atom representations, our experiments are restricted to $\alpha$-carbon atoms to facilitate direct comparisons.
@@
[…]
FPocket [Le Guilloux et al., 2009] $\times$ 31.5 23.2 34.1 25.4 30.0
SiteHound [Hernandez et al., 2009] $\times$ 36.4 23.1 38.9 29.4 34.9
MetaPocket2 [Macari et al., 2019] $\times$ 37.3 25.8 37.5 32.8 37.7
DeepSite [Jiménez et al., 2017] $\times$ 34.0 29.1 37.4 27.4 33.9
P2Rank [Krivák and Hoksza, 2018] $\times$ 49.8 38.6 47.4 56.5 45.3
ESM2_150M [Lin et al., 2023b] $\surd$ 19.6 16.6 20.5 18.9 22.0
GearNet [Zhang et al., 2022b] $\surd$ 39.9 35.8 41.0 36.4 41.3
Siamdiff [Zhang et al., 2023d] $\surd$ 37.7 31.0 40.7 35.3 40.3
Vabs-Net [Zhao et al., 2024] $\surd$ - - - - 56.3
@@
[…]
E3Bind [Zhang et al., 2022a] / 25.6 7.2 / / /
KarmaDock [Zhang et al., 2023c] / 56.2 / / / /
DiffDock(Pocket) [Corso et al., ] / 51.8 2.0 / 60.7 1.9
DiffDock [Corso et al., ] 15.15 36.09 3.35 21.76 43.52 2.46
[…]
C.5 Molecular Docking
@@
We evaluate Uni-3DAR against 13 baselines, including five classical docking software tools and eight deep learning–based methods. Most existing deep learning approaches rely on complex featurizations, such as using protein language model embeddings (e.g., from ESM2 [Lin et al., 2022]). To simplify and unify the molecular tasks, we omit these complicated features in Uni-3DAR and instead use only atom types and coordinates. We also adopt a full-atom representation of the protein pocket to enhance expressive power. We frame docking as an autoregressive generation task by embedding both the pocket and the RDKit conformation as two frames, concatenating them into a single input sequence, and training the model to generate the docked molecule conformation as a new frame sequence. For further simplicity, we do not impose constraints such as matching the number and types of atoms in the output frame to those of the input molecule. Also, we do not train a separate scoring model for pose ranking. Instead, we use the cumulative probability derived from autoregressive generation to score each generated pose. We train Uni-3DAR for 300k steps (approximately 300 epochs) with a batch size of 16. The learning rate schedule follows the same configuration as the experiments detailed in section˜3.1. The training is completed in approximately one day on 4 NVIDIA A100 GPUs.
@@ Results
Experimental results are summarized in Table˜7. Uni-3DAR outperforms to the state-of-the-art method, SurfDock, demonstrating similar percentages of poses with RMSD below 1Å and 2Å. Notably, Uni-3DAR excels in generating higher-quality poses, reflected by its lower median RMSD values. However, Uni-3DAR exhibits slightly inferior performance in selecting Top-5 poses for challenging cases, as evidenced by a lower percentage of poses with RMSD below 2Å in the Top5-RMSD evaluation (72.38% vs. 75.11% for SurfDock). This gap may arise because the scoring module in Uni-3DAR has not been explicitly trained and is only exposed to ground-truth conformations during the training phase. Addressing this limitation by training a dedicated scoring module could potentially enhance its selection performance. Moreover, since Uni-3DAR avoids complex feature engineering, its docking accuracy might further benefit from multitask learning strategies, emphasizing the promise of a unified foundational model for molecular applications.
C.6 Molecular Property Prediction via Pretraining
@@
We utilize the same pretraining dataset as employed by Uni-Mol [Zhou et al., 2023b] and SpaceFormer [Lu et al., 2025], comprising approximately 19 million molecules. For downstream evaluations, we follow the datasets and evaluation settings used by the state-of-the-art SpaceFormer [Lu et al., 2025]. These include a 20K dataset predicting electronic properties (HOMO, LUMO, GAP), a 21K dataset targeting energy properties (E1-CC2, E2-CC2, f1-CC2, f2-CC2), and an 8K dataset predicting mechanical and electronic properties (Dipmom, aIP, and D3 Dispersion Corrections). Data splits align exactly with SpaceFormer’s methodology [Lu et al., 2025]. Performance across all tasks is measured using the Mean Absolute Error (MAE) metric.
@@
Our baselines encompass several prominent models, including Uni-Mol [Zhou et al., 2023b], Mol-AE [Yang et al., 2024], 3D Infomax [Stärk et al., 2022a], GROVER [Rong et al., 2020], GEM [Fang et al., 2022], and the most recent state-of-the-art method, SpaceFormer [Lu et al., 2025]. For pretraining, we use the proposed masked next-token prediction as pretraining task, training the model for 500k steps with a batch size of 128. The peak learning rate is set to 3e-4, incorporating a 10% linear warmup followed by cosine decay, requiring approximately 11.5 hours on 8 NVIDIA 4090 GPUs.
@@
LUMO $\downarrow$
GAP $\downarrow$
E1-CC2 $\downarrow$
E2-CC2 $\downarrow$
f1-CC2 $\downarrow$
f2-CC2 $\downarrow$
Dipmom $\downarrow$
[…]
_corr (eV)
GROVER [Rong et al., 2020]
0.0075
$\pm$ 2.0e-4
0.0086
$\pm$ 8.0e-4
0.0109
$\pm$ 1.4e-3
0.0101
$\pm$ 9.7e-4
0.0129
$\pm$ 4.6e-4
0.0219
$\pm$ 3.5e-4
0.0401
$\pm$ 1.2e-3
0.0752
$\pm$ 1.1e-3
0.1467
$\pm$ 1.5e-2
0.2516
$\pm$ 5.3e-2
GEM [Fang et al., 2022]
0.0068
$\pm$ 7.0e-5
0.0080
$\pm$ 2.0e-5
0.0107
$\pm$ 1.9e-4
0.0090
$\pm$ 1.3e-4
0.0102
$\pm$ 2.3e-4
0.0170
$\pm$ 4.3e-4
0.0352
$\pm$ 5.4e-4
0.0289
$\pm$ 1.2e-3
0.0207
$\pm$ 2.6e-4
0.0077
$\pm$ 6.6e-4
3D Infomax [Stärk et al., 2022a]
0.0065
$\pm$ 1.0e-5
0.0070
$\pm$ 1.0e-4
0.0095
$\pm$ 1.0e-4
0.0089
$\pm$ 2.0e-4
0.0091
$\pm$ 3.0e-4
0.0172
$\pm$ 4.0e-4
0.0364
$\pm$ 9.0e-4
0.0291
$\pm$ 1.7e-3
0.0526
$\pm$ 1.4e-4
0.2285
$\pm$ 7.5e-3
Uni-Mol [Zhou et al., 2023b]
0.0052
$\pm$ 2.0e-5
0.0060
$\pm$ 6.0e-5
0.0081
$\pm$ 4.0e-5
0.0067
$\pm$ 4.0e-5
0.0080
$\pm$ 4.0e-5
0.0143
$\pm$ 2.0e-4
0.0309
$\pm$ 9.4e-4
0.0106
$\pm$ 3.1e-4
0.0095
$\pm$ 6.4e-4
0.0047
$\pm$ 5.6e-4
Mol-AE [Yang et al., 2024]
0.0050
$\pm$ 8.0e-5
0.0057
$\pm$ 4.7e-4
0.0080
$\pm$ 8.0e-5
0.0070
$\pm$ 6.0e-5
0.0080
$\pm$ 4.0e-5
0.0140
$\pm$ 4.0e-5
0.0307
$\pm$ 1.3e-3
0.0113
$\pm$ 4.7e-4
0.0103
$\pm$ 1.3e-4
0.0077
$\pm$ 1.3e-3
SpaceFormer [Lu et al., 2025]
0.0042
$\pm$ 1.0e-5
0.0040
$\pm$ 2.0e-5
0.0064
$\pm$ 1.2e-4
0.0058
$\pm$ 8.0e-5
0.0074
$\pm$ 8.4e-5
0.0142
$\pm$ 3.7e-4
0.0294
$\pm$ 7.1e-4
0.0083
$\pm$ 5.0e-4
0.0090
$\pm$ 5.9e-4
0.0053
$\pm$ 1.2e-3
Uni-3DAR
0.0048
$\pm$ 2.1e-5
0.0044
$\pm$ 3.2e-5
0.0065
$\pm$ 8.8e-5
0.0056
$\pm$ 2.2e-5
0.0067
$\pm$ 2.0e-5
0.0134
$\pm$ 7.0e-5
0.0286
$\pm$ 1.6e-4
0.0114
$\pm$ 6.9e-4
0.0127
$\pm$ 1.1e-4
0.0052
$\pm$ 3.2e-4
C.7 Polymer Property Prediction via Pretraining